Can someone help me solve this by using the sqare root property?
(x+8)^2=81

Answer by usyim88hk(158)

(x+8)^2 = 81
*square root both sides
(x+8) = 9 OR (x+8) = -9
**For (x+8) = 9
*subtract 8 from both sides
x = 1
**For (x+8) = -9
*subtract 8 from both sides
x = -17
So there are two answers for this equation
x = 1 or -17